Maria drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 9 hours. When Maria drove home, there was no traffic and the trip only took 4 hours. If her average rate was 40 miles per hour faster on the trip home, how far away does Maria live from the mountains?
Do not do any rounding.

Answer by mananth(16949)

let Maria live x miles from the mountains
Let her rate to mountains be y
t=d/r
x/y = 9
x=9y
Return trip 4 hours
rate = y+40
x/(y+40) =4
substitute x
9y/(y+40)=4
9y=4y+160
9y-4y=160
5y=160
y= 160/5
y=32
x=9y
x=9*32= 288 miles
distance to the mountain = 288 miles